OpenStudy (anonymous):

Write x^2 - 2x - 3 = 0 in the form (x - a)^2 = b, where a and b are integers. (x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

No I don't.

OpenStudy (anonymous):

Okay, cool :) I will help

OpenStudy (anonymous):

Thanks. I have been in this exam for coming up on 5 hours. Please bare with me.

OpenStudy (anonymous):

We have: x^2 - 2x - 3 = 0 To complete the square, we take the middle number, halve it and then square it. (-2 / 2)^2 = 1 Now we have the answer, '1', we add this and take it to the end of the equation: x^2 - 2x + 1 - 1 - 3 = 0 This then allows us to simplify the first part as a perfect square: x^2 - 2x + 1 - 1 - 3 = 0 (x - 1)^2 - 1 - 3 = 0 (x - 1)^2 = 4 So the answer is: (x - 1)^2 = 4

OpenStudy (anonymous):

Thank you SO much @Riddellikins

OpenStudy (anonymous):

You're welcome :)